Present Value of Bonds Payable; Premium Moss Co. issued $105,000 of five-year, 12% bonds with interest payable semiannually, at a market (effective) interest rate of 9% Determine the present value of the bonds payable, using the present value tables in Exhibit 8 and Exhibit 10 Note: Round to the nearest dollar. Previous Next
Answer #1
Present Value of bonds payable $       1,17,463
Working:
# 1 Semi annual coupon payment = Par Value * Semi annual coupon rate
= 105000*12%*6/12
= $             6,300
# 2 Present Value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.045)^-10)/0.045 i 9%*6/12 = 0.045
=             7.91272 n 5*12/6 = 10
# 3 Present Value of 1 = (1+i)^-n
= (1+0.045)^-10
=             0.64393
# 4 Present value of coupon payment $             6,300 *    7.91272 = $           49,850
Present Value of Par Value $       1,05,000 *    0.64393 = $           67,612
Total Present Value of Cash flows $       1,17,463